c. P=50 - 0.1Q and P=$20

Question is how can we know that Q=500-10P?

Could you pls be more specific?


1
Expert's answer
2021-01-07T08:27:38-0500

"Q = 500-10P"

"P = 50 - 0.1Q"

"P = $ 20"

"P = 50 - 0.1 Q"

"20 = 50 - 0.1Q"

"0.1Q = 50 - 20"

"Q = 30\/0.1"

"Q = 300"

"Q = 500 - 10 P"

"300 = 500 - 20(10)"

"300 = 500 - 200"

"300 = 300"

"At equilibrium Q = P"